# Quantifying Privacy Leakage in Split Inference via Fisher-Approximated Shannon Information Analysis

### v1

@@ 5.3.1 Intrinsic Data Characteristics
Figure 8: The impact of dataset complexity on privacy leakage. The left diagram is the FSInfo among different split layers of VGG-5 on MNIST and CIFAR-10. The right diagram is the raw inputs (Ref.) and the reconstructed inputs from split point 2 (Inv.(2)).
CIFAR-10 comprises diverse RGB images across 10 classes, whereas MNIST contains grayscale handwritten digits. Normalized to $[-1,1]$, CIFAR-10 exhibits richer pixel value diversity compared to MNIST’s bimodal distribution peaking at -1 and 1 (Appendix B). To investigate data characteristics’ impact on privacy leakage, we analyze FSInfo for both datasets on VGG-5 across four split points (Figure 8, left). Results demonstrate consistently lower privacy leakage for CIFAR-10, aligning with Liu et al.’s findings (Liu et al., 2022) that simpler images facilitate reconstruction. We attribute this discrepancy to dataset complexity and value space dimensions. Visual comparisons of raw and reconstructed images (Figure 8, right) reveal that MNIST adversaries primarily handle binary color patterns, while CIFAR-10 reconstruction requires capturing complex color distributions, suggesting that richer information content increases reconstruction difficulty.
5.3.2 Model Width and Depth
@@
Figure 9: The impact of model size on privacy leakage on different models. Reconstructed inputs of the raw input (left) along with their FSInfo are displayed. The ResNet-18 architecture is displayed for three different base block channel sizes.
@@
The model size (number of layers and the output size of a layer) influences neural network privacy characteristics. Figure 9 demonstrates that both reconstruction quality and FSInfo increase with channel size in ResNet-18, as wider networks exhibit more pronounced neuron activation patterns. Interestingly, the deeper network (ResNet-34) may not always decrease the privacy leakage, because after being trained over 20 epochs, the ResNet-34 has a higher test accuracy than ResNet-18 on CIFAR-10, which may lead to a higher privacy leakage (see Section 5.3.3). Jiang et al. (Jiang et al., 2022) suggest that additional layers hinder leakage through increased transformations, which is consistent with the results of Figure 10 when SP is from 5 to 6. However, Figure 10 reveals this intuition may be incomplete, as later split points (e.g., SP is from 3 to 4) can exhibit greater leakage with increased width. Notably, at SP=8, MNIST shows lower leakage than CIFAR-10 due to MaxPooling’s differential impact on feature map dimensions (7x7→3x3 vs. 8x8→4x4). Overall, network width substantially amplifies privacy leakage, while the increased depth degrades privacy leakage more moderately.
5.3.3 Overfitting
@@
Following established methodologies (Liu et al., 2022; Song & Mittal, 2021; Carlini et al., 2019), we refer to the training epochs as the overfitting level. Figure 11 illustrates the relationship between epoch count and privacy leakage (FSInfo) for both VGG-5 and VGG-9, revealing a characteristic U-shaped pattern. Initially, untrained models exhibit high sensitivity to input stimuli, with a substantial gradient scale of smashed data w.r.t. the raw input responses leading to significant information leakage. As training progresses, this sensitivity gradually diminishes. However, prolonged training enables model parameters to memorize training data information, which subsequently manifests in smashed data during consequent testing (Achille et al., 2019), explaining the observed first-decline-then-rise FSInfo trajectory.
6 Conclusion
@@
We correlate the privacy leakage with the DRA adversary’s certainty and formulate it with conditional entropy, which can lower bound the adversary’s reconstruction error in both average- and worst-case. We obtain an operationally computable privacy metric, FSInfo, by transforming Fisher information into Shannon quantity. Experimental results showed that FSInfo aligns well with the human visual judgments on privacy leakage and the objective attack accuracy metrics. Also, we experimentally analyze factors that affect privacy leakage. Future work may extend the analysis of privacy leakage to more complex models (e.g., generative models) and other exposed information carriers (e.g., gradients in SL). Besides, the analysis of the privacy leakage determinants is conducted experimentally; future works may study theoretically and derive more interpretable results. Also, we experimentally analyze the determinants of privacy leakage.
Appendix A More Related Work
A.1 Data Reconstruction Attacks to SI
@@
Discriminative Reconstruction Attacks (DRAs) employ two primary approaches: neural network (NN)-based methods, which learn mappings from observable outputs to raw inputs, and maximum likelihood estimation (MLE)-based methods, utilizing optimization techniques like stochastic gradient descent (He et al., 2019). In federated learning, where only aggregated parameters and gradients are observable, MLE-based methods dominate, typically optimizing gradient matching objectives (euclidean or cosine similarity) (Zhang et al., 2023; Geiping et al., 2020; Jin et al., 2021).
SI scenarios favor NN-based methods due to the direct one-to-one correspondence between raw inputs and smashed data, providing ample training pairs for inverse neural networks (Yin et al., 2023; Pasquini et al., 2021). Recent SI-focused DRA research primarily enhances attack accuracy under constrained conditions, such as lacking auxiliary datasets (Yin et al., 2023) or facing defense mechanisms (Yang et al., 2022). Empirical evidence demonstrates superior reconstruction quality of NN-based methods compared to MLE-based approaches (Yin et al., 2023).
A.2 Information in Deep Neural Network
@@
In the information theory, entropy $H(X)$ is a measure of the uncertainty of a random variable $X$ and (Shannon) mutual information $I(X;Y)$ between random variables $X$ and $Y$ is the reduction in the uncertainty of $X$ due to the knowledge of $Y$, and vice versa (Cover, 1999):
$I(X;Y)=H(X)-H(X|Y)=H(Y)-H(Y|X).$ (24)
@@
Previous studies (Tishby & Zaslavsky, 2015; Yu et al., 2020) analyze DNNs by seeing the inputs, predictions, and outputs of one hidden layer (representations) as random variables $X$, $Y$, and $Z$. Model optimization can be done by minimizing $I(X;Z)$ and maximizing $I(Z;Y)$ under the information bottleneck (IB) tradeoff (Shwartz-Ziv & Tishby, 2017; Yu et al., 2020).
Appendix B Experimental Implementations
B.1 Implementation
@@
Our experiments were performed on a server equipped with two NVIDIA Geforce RTX 4090 24GB GPUs, two Intel Xeon Silver 4310 CPUs, and 256GB of memory. We implemented the proposed metric with Python 3.8.13, and Pytorch 1.12.1.
B.2 Datasets
Table 2: Statistics of Datasets
Dataset # Train # Test # Feature # Class
CIFAR-10 50,000 10,000 32x32x3 10
MNIST 60,000 10,000 28x28x1 10
@@ B.2 Datasets
We evaluated our metrics on four tasks with two real-world datasets (CIFAR-10 (Krizhevsky et al., 2009) and MNIST (LeCun et al., 1998)) as shown in Table 2. Both datasets are normalized into $[-1,1]$.
CIFAR10 (Krizhevsky et al., 2009): Comprising 50,000 training samples and 10,000 test samples across 10 classes, CIFAR-10 dataset is a benchmark in computer vision with categories like airplanes, automobiles, and birds. Each image is a 32x32 pixel color image with three channels.
MNIST (LeCun et al., 1998): Comprising 60,000 training samples and 10,000 test samples across 10 digits, with each image being a 28x28 gray-scale image.
Figure 12 plots the distribution of the first image of CIFAR-10 and MNIST respectively. We bin the continuous pixel value of two images into 100 (for CIFAR-10) and 10 (for MNIST) discrete intervals between -1 and 1.
B.3 Models
Table 3: Statistics of Models. Convolution layers are denoted by C, followed by the number of filters; MaxPooling, AveragePooling, and Adaptive AveragePooling layers are MP, AP, and AAP; Fully Connected layer is FC with the number of neurons; Basic Block of ResNet is BB with the channel size.
@@ B.3 Models
We use four models including VGG-5 (Simonyan & Zisserman, 2014), VGG-9 (Simonyan & Zisserman, 2014), ResNet-18 (He et al., 2016), ResNet-34 (He et al., 2016), as shown in Table. 3.
@@
The number of critical layers and architecture of these models implemented for CIFAR-10 are presented in Table 3. We also identify the default split layer in the brackets. All models are trained until the test accuracy tends to stabilize by default (approximately 20 epochs). The Gaussian distribution for initialize the parameters of $f_{\theta 1}$ is $\mathcal{N}(0,0.01)$.
Appendix C Details of Compared Methods and Metrics
C.1 Compared Methods
Diagonal Fisher information leakage (dFIL) (Guo et al., 2022; Maeng et al., 2023) views the DRA as parameter estimation and utilizes Fisher information to bound the estimation error as follows:
$\mathbb{E}[\|\hat{x}-x\|_{2}^{2}/d]\geq\frac{1}{dFIL}$ (25)
$dFIL=Tr(\mathcal{I}_{z}(x))/d,$ (26)
where the dimension of input $x$ is $d$, $Tr(\cdot)$ denotes the matrix trace, $\hat{x}$ is the reconstructed input, and $\mathcal{I}(\cdot)$ is the Fisher information matrix of smashed data $z$ with respect to $x$. Following Maeng et al.(Maeng et al., 2023), we use the reciprocal of dFIL (1/dFIL), the high value of which indicates less privacy leakage.
DLoss (Deng et al., 2024; Székely et al., 2007; Sun et al., 2021, 2022) calculates the privacy leakage as the distance correlation between raw input $X$ and smashed data $Z$, which can measure the nonlinear dependence between these random vectors:
$\text{DLoss}(X,Z)=\sqrt{\mathcal{R}^{2}(X,Z)},$ (27)
[…]
where the $\mathcal{R}^{2}(X,Z)$ is the distance covariance between $X$ and $Z$.
Mutual Information (MI) (Mireshghallah et al., 2020) calculates the privacy leakage as the Shannon mutual information using the Information Theoretical Estimators Toolbox (Ghosh et al., 2021) as:
[…]
where $x$ and $z$ are the raw input and smashed data.
C.2 Metrics
Visual invertibility measures the similarity of the raw input and the reconstructed one by human perception (Sun et al., 2024). We visualize the reconstructed images and observe their clarity and similarity to the raw input. This metric offers semantic judgment for privacy information observed by the adversary. A higher visual quality means a larger privacy leakage.
SSIM is an objective criterion for measuring the pixel-wise similarity of two images with values ranging from 0 to 1. A higher SSIM means a larger privacy leakage.
MSE is also an objective criterion, which measures the error or distance between two images. A large MSE indicates less privacy leakage.
Appendix D Examining Defense mechanisms
Figure 13: Changing trend of FSInfo and MSE with the growth of Gaussian noise scale on VGG-5.
(a) VGG-5
(b) VGG-9
(c) Resnet-18
Figure 14: FSInfo under Nopeek defense mechanisms on CIFAR-10.
As mentioned in Section 4.3, as a privacy metric, FSInfo can be used for analyzing and supervising the privacy risks of DNN-based products. In this section, we experimentally validate that, practitioners can also use FSInfo to assess the split inference system employed defense mechanisms, demonstrating the practical relevance of our study.
The unique characteristics of split inference necessitate distinct privacy protection approaches compared to traditional federated learning. In split inference, the smashed data maintains a one-to-one correspondence with the raw input, rather than representing an aggregated form of the raw input. Consequently, the batch size, which often serves as a privacy consideration in federated learning, may not significantly affect privacy leakage in the smashed data of split inference (Geiping et al., 2020). This characteristic also explains why standard DP is unsuitable for split learning scenarios—the standard DP definition of neighboring databases requires that outputs for any two samples remain nearly indistinguishable, regardless of their differences (Singh et al., 2023). Consequently, obfuscation-based defense mechanisms, particularly those utilizing direct noise injection or adversarial representation learning techniques, have become predominant in split inference systems (Singh et al., 2024).
We apply the proposed FSInfo framework to split inference systems incorporating these two obfuscation defense mechanisms:
  • Gaussian noise (Singh et al., 2024) directly adds noise from Gaussian distribution on the raw smashed data before sending it to the server.
  • Nopeek (Vepakomma et al., 2020) uses adversarial representation learning methods in the model training phase to guide the bottom model to output smashed data with less information of raw input. For raw input data $X$, smashed data $Z$, true labels $Y_{true}$ and predicted labels $Y$, the loss function of Nopeek is:
    $\alpha_{1}DLoss(X,Z)+\alpha_{2}CCE(Y_{true},Y),$
    where the $DLoss$ is the privacy metric metiond in Section 5.1.2, and the $CCE$ is the cross-entropy loss function.
Figure 13 shows the changing trend of FSInfo and MSE with the growth of Gaussian noise scale on VGG-5 (with the default split layer mentioned in Section 6 in the paper) and CIFAR-10, where we varied the Gaussian noise from 0.1 to 50. As we can tell, FSInfo thinks an SL system with higher defense strength (higher noise scale) may leak less leakage.
We train the VGG-5, VGG-9, and ResNet-18 with Nopeek on CIFAR-10 to get the protected model, where both $\alpha_{1}$ and $\alpha_{2}$ are 0.5. Figure 14 shows the FSInfo on different split points on both raw and protected models. The results demonstrate that after defense, the smashed data leaks less information about the raw input. Also, Figure 14 (a) shows that privacy leakage of the 5-th split point (the first fully connected layer of VGG-5) of the protected model is larger than that of the 4-th, indicating that Nopeek may be less effective for defending privacy leakage from the first fully connected layer after the convolutional layers.
Appendix E Proof of Theorem 4.2

### v2

@@ V-F The Impact of Intrinsic Data Characteristics
Figure 8: The impact of dataset complexity on privacy leakage. The left diagram is the FSInfo among different split layers of VGG-5. The right diagram is the raw inputs (Ref.) and the reconstructed inputs from split point 2 (Inv.(2)).
@@
Figure 9: The impact of model size on privacy leakage on different models. Reconstructed inputs, raw input (Ref.), and FSInfo are displayed. The ResNet-18 architecture is displayed for three different base block channel sizes.
@@
CIFAR-10 comprises diverse RGB images across 10 classes, whereas MNIST contains grayscale handwritten digits. Normalized to [1,1][-1,1], CIFAR-10 exhibits richer pixel value diversity compared to MNIST’s bimodal distribution peaking at -1 and 1 (see Appendix). To investigate data characteristics’ impact on privacy leakage, we analyze FSInfo for both datasets across four split points (Fig. 8, left). Results demonstrate consistently lower privacy leakage for CIFAR-10, aligning with Liu et al.’s findings [liu2022ml] that simpler images facilitate reconstruction. We attribute this discrepancy to dataset complexity and value space dimensions. Visual comparisons of raw and reconstructed images (Fig. 8, right) reveal that MNIST adversaries primarily handle binary color patterns, while CIFAR-10 reconstruction requires capturing complex color distributions, suggesting that richer information content increases reconstruction difficulty.
V-G The Impact of Model Width and Depth
The model size (number of layers and the output size of a layer) influences neural network privacy characteristics. Fig. 9 demonstrates that both reconstruction quality and FSInfo increase with channel size in ResNet-18, as wider networks exhibit more pronounced neuron activation patterns. Interestingly, the deeper network (ResNet-34) may not always decrease the privacy leakage, because after being trained over 20 epochs, the ResNet-34 has a higher test accuracy than ResNet-18 on CIFAR-10, which may lead to a higher privacy leakage (see Section V-H). Jiang et al. [jiang2022fedsyl] suggest that additional layers hinder leakage through increased transformations, which is consistent with the results of Fig. 10 when SP is from 5 to 6. However, Fig. 10 reveals this intuition may be incomplete, as later split points (e.g., SP is from 3 to 4) can exhibit greater leakage with increased width. Notably, at SP=8, MNIST shows lower leakage than CIFAR-10 due to MaxPooling’s differential impact on feature map dimensions (7x7→3x3 vs. 8x8→4x4). Overall, network width substantially amplifies privacy leakage, while the increased depth degrades privacy leakage more moderately.
V-H The Impact of Overfitting
@@
Following established methodologies [liu2022ml, sec21-evaluation-ml, Sec19-secret-share], we refer to the training epochs as the overfitting level. Fig. 11 illustrates the relationship between epoch count and privacy leakage (FSInfo) for both VGG-5 and VGG-9, revealing a characteristic U-shaped pattern. Initially, untrained models exhibit high sensitivity to input stimuli, with a substantial gradient scale of smashed data w.r.t. the raw input responses leading to significant information leakage. As training progresses, this sensitivity gradually diminishes. However, prolonged training enables model parameters to memorize training data information, which subsequently manifests in smashed data during consequent testing [achille2019information], explaining the observed first-decline-then-rise FSInfo trajectory.
VI Conclusion
@@
We correlate the privacy leakage with the DRA adversary’s certainty and formulate it with conditional entropy, which can lower bound the adversary’s reconstruction error in both average- and worst-case. We obtain an operationally computable privacy metric, FSInfo, by transforming Fisher information into Shannon quantity, and we design a defense method, FSInfoGuard, with a strong utility-privacy trade-off. Experimental results show that FSInfo reliably tracks actual privacy leakage across models, datasets, and defense strengths, exhibiting strong consistency with empirical attack performance and even human-perceived reconstruction quality. Also, FSInfo can be used to guide the design of defense mechanisms and to provide insights into key system components. Future work may extend the analysis of privacy leakage to other exposed information carriers (e.g., gradients in SL). Besides, the analysis of the privacy leakage determinants is conducted experimentally; future works may study it theoretically and derive more interpretable results.
More Related Work
-A Data Reconstruction Attacks to SI
@@
Data Reconstruction Attacks (DRAs) employ two primary approaches: neural network (NN)-based methods, which learn mappings from observable outputs to raw inputs, and maximum likelihood estimation (MLE)-based methods, utilizing optimization techniques like stochastic gradient descent [blackbox]. In federated learning, where only aggregated parameters and gradients are observable, MLE-based methods dominate, typically optimizing gradient matching objectives (Euclidean or cosine similarity) [zhang2023survey, geiping2020inverting, jin2021cafe].
SI scenarios favor NN-based methods due to the direct one-to-one correspondence between raw inputs and smashed data, providing ample training pairs for inverse neural networks to achieve a better attack performance than the MLE-based [yin2023ginver, pasquini2021unleashing]. Recent methods in this line of strategy have focused on adjusting the training set, loss function, and decoder network architecture to prevent performance degradation in more practical scenarios. For example, the adversary does not have auxiliary datasets [yin2023ginver] or face defense mechanisms [yang2022measuring]. We aim to evaluate the system’s robustness and provide guidelines for defense. In this case, we set loose conditions for the adversaries and allow them to use the basic but well-performing inverse network method [blackbox] in our experiments.
-B Information in Deep Neural Network
@@
In the information theory, entropy $H(X)$ is a measure of the uncertainty of a random variable $X$ and (Shannon) mutual information $I(X;Y)$ between random variables $X$ and $Y$ is the reduction in the uncertainty of $X$ due to the knowledge of $Y$, and vice versa [cover1999elements]:
I(X;Y)=H(X)H(X|Y)=H(Y)H(Y|X).I(X;Y)=H(X)-H(X|Y)=H(Y)-H(Y|X). (15)
@@
Previous studies [tishby2015deep, yu2020understanding] analyze DNNs by seeing the inputs, predictions, and outputs of one hidden layer (representations) as random variables $X$, $Y$, and $Z$. Model optimization can be done by minimizing $I(X;Z)$ and maximizing $I(Z;Y)$ under the information bottleneck (IB) tradeoff [shwartz2017opening, yu2020understanding].
Experimental Implementations
-C Implementation
@@
Our experiments were performed on a server equipped with two NVIDIA GeForce RTX 4090 24GB GPUs, two Intel Xeon Silver 4310 CPUs, 256GB of memory, and Debian 6.1.119-1. We implemented the proposed metric with Python 3.10 and Pytorch 2.4.1.
-D Datasets
TABLE I: Statistics of Datasets
Category Dataset # Train # Test # Feature # Class
Tabular Purchase100 158k 39k 600 100
Home Credit 246k 62k 121 2
Image CIFAR-10 50k 10k 32x32x3 10
MNIST 60k 10k 28x28x1 10
CelebA 163K 20k 178x218x3 2
TinyImageNet 100k 10k 64x64x3 4
Text GLUE/SST-2 67k 1.8k - 2
@@
We evaluated our metrics on seven real-world datasets, covering tabular, text, and image datasets, as shown in Table I. All image data are first rescaled to [0,1] and then normalized to [-1,1] with a variance and mean of 0.5. For two tabular datasets, we one-hot encode the discrete columns and rescale the continuous columns to [-1,1]. For CelebA [liu2015faceattributes], we built the binary attractiveness classification task following [nips2023gan] and resize the image to 112x112x3. For TinyImageNet, we built 4-class classification task follow [sec-xu2025risk].
CIFAR10 [krizhevsky2009learning]: Comprising 50,000 training samples and 10,000 test samples across 10 classes, CIFAR-10 dataset is a benchmark in computer vision with categories like airplanes, automobiles, and birds. Each image is a 32x32 pixel color image with three channels.
MNIST [lecun1998gradient]: Comprising 60,000 training samples and 10,000 test samples across 10 digits, with each image being a 28x28 gray-scale image. Fig. 12 plots the distribution of the first image of CIFAR-10 and MNIST, respectively. We bin the continuous pixel value of two images into 100 (for CIFAR-10) and 10 (for MNIST) discrete intervals between -1 and 1. We can see that the probability distribution of pixel values in MNIST is more concentrated around specific values, while CIFAR-10 shows a more uniform spread across the range.
CelebA [liu2015faceattributes]: is a large-scale facial attributes dataset containing over 200,000 celebrity images, which are divided into training set 162770 (proportion 0.8), verification set 19867 (proportion 0.098), and test set 19962 (proportion 0.096). Each image has 40 binary attribute annotations and 5 landmark locations. It is widely used as a benchmark for tasks such as face recognition, facial attribute prediction, and generative adversarial networks (GANs) due to its large diversity, rich annotations, and high quality.
TinyImageNet [le2015tiny]: is a downsized and more computationally manageable variant of the ImageNet Large Scale Visual Recognition Challenge (ILSVRC). It contains 100,000 images across 200 object classes, with each class having 500 training, 50 validation, and 50 test images.
Purchase100 (Purchase) [sec21-evaluation-ml]: is based on Kaggle’s Acquire Valued Shoppers Challenges, which has 600 binary features indicating whether a good has been purchased or not, for identifying different shopping styles of a customer. We use the simplified version proposed by [sec21-evaluation-ml].
Home Credit (Credit) [home-credit-kaggle]: is based on Kaggle’s Home Credit Default Risk Challenge, to predict the credit applicant’s default likelihood.
GLUE/SST-2 [wang2018glue]: GLUE is a multi-task benchmark for natural language understanding, where the SST-2 is designed for sentiment analysis.
-E Models
TABLE II: Statistics of Models. Convolution layers are denoted by C, followed by the number of filters; MaxPooling, AveragePooling, and Adaptive AveragePooling layers are MP, AP, and AAP; Fully Connected layer is FC with the number of neurons; Basic Block of ResNet is $B$ with the channel size. EMB is the embedding layer of the transformer, and Tr is the transformer block.
@@
FCN for Purchase 5 FC1024-FC512(default)-FC256-FC128-FC100
FCN for Credit 4 FC512-FC128(default)-FC32-FC1
DistilBert for GLUE/SST-2 8 EMB768-Tr768(defulat)-[Tr768]*4-FC768-FC2
We use seven models, including VGG-5 [simonyan2014very], VGG-9 [simonyan2014very], ResNet-18 [he2016deep], ResNet-34 [he2016deep], a 5-layer linear model for Purchase, a 4-layer linear model for Credit, and DistilBert [sanh2019distilbert], as shown in Table. II. For image data, we only give the model architecture for CIFAR-10 as a representative case, since the architectures for other image datasets involve only minor modifications on the input and output layers.
@@
The number of critical layers and architecture of these models are presented in Table II. We also identify the default split layer in the brackets. All models except for DistilBert are trained until the test accuracy tends to stabilize by default (approximately 20 epochs). For DistilBert, we use the pretrained weights from HuggingFace.
Detailed Implementations for FSInfo calculation
We give the detailed implementation of FSInfo in Table III.
TABLE III: Numerical calculation procedure for FSInfo.
Step Actions
1 Calculate the $d_{x}log(2\pi e)$.
2 Calculate the diagonal Fisher information Matrix (FIM).
2.1 For the small models (VGG-5, VGG-9, ResNet-18, ResNet-34, and MLPs for Credit and Purchase), we use intuitive and simple implementations:
2.1.1 Calculate the Jacobian of $z$ w.r.t. $x$, $J_{f_{\theta 1}(x)}$ , using ”torch.autograd.functional.jacobian”.
2.1.2 Calculate the Fisher Information Matrix $F_{z|x}=\frac{1}{\sigma^{2}}J_{f_{\theta 1}(x)}^{\top}J_{f_{\theta 1}(x)}$ and get the diagonal vectors $\Lambda=(\lambda_{1},\dots,\lambda_{d_{x}})$.
2.2 For big models (DistilBert), we want to save the memory cost:
2.2.1 Only calculate the diagonal item of $J_{f_{\theta 1}(x)}^{\top}J_{f_{\theta 1}(x)}$ instead of calculating the whole matrix by using ”torch.autograd.functional.jvp” for memory efficiency. Then we get the diagonal vectors of $J_{f_{\theta 1}(x)}^{\top}J_{f_{\theta 1}(x)}$.
2.2.2 Multiply each diagonal item of $J_{f_{\theta 1}(x)}^{\top}J_{f_{\theta 1}(x)}$ by $\frac{1}{\sigma^{2}}$. Then we get the diagonal vectors $\Lambda=(\lambda_{1},\dots,\lambda_{d_{x}})$.
3 Calculate the term $\log\det(F_{z|x})=\Sigma_{i}^{d_{x}}\log(\lambda_{i}+1*e^{-10)})$.
4 Get the FSInfo for $x$ FSInfo = $\frac{1}{2d_{x}}[d_{x}log(2\pi e)-\log\det(F_{z|x})]$.
Proof of Theorem 4.2